# Intermediate scattering function of a gravitactic circle swimmer

### v1

@@ III Derivation of the intermediate scattering function
Besides the equations of motion for the individual trajectories from Sec. II, dynamical properties are equivalently encoded in the propagator :-$\mathbb{P}\coloneq\mathbb{P}(\mathbf{r},\vartheta,t|\vartheta_{0})$, which is defined as the conditional probability distribution of the particle displaced by $\mathbf{r}$ at lag time $t$ with orientation $\vartheta$, given it started with orientation $\vartheta_{0}$. By standard means Risken (1996), the associated Fokker-Planck equation can be derived
@@
It is favorable to solve directly for the spatial Fourier transform of the propagator ~:-~$\tilde{\mathbb{P}}\coloneq\tilde{\mathbb{P}}(\mathbf{k},\vartheta,t|\vartheta_{0})$
~$\tilde{\mathbb{P}}=\int_{\mathbb{R}^{2}}\!\mathrm{d}\,\textbf{r}\exp(-\textsf{i}\textbf{k}\cdot\textbf{r})\mathbb{P}(\mathbf{r},\vartheta,t|\vartheta_{0}).$ (5)
@@
-:$\eqcolon(\mathcal{L}+\delta\mathcal{L}_{\mathbf{k}})\tilde{\mathbb{P}},-$ (6)
where the operator $\mathcal{L}$ encodes the angle-dependent motion containing the drifts $\omega$ and $\gamma$, and the rotational diffusion. The $\mathbf{k}$-dependent operator $\delta\mathcal{L}_{\mathbf{k}}$ encodes the deterministic active motion for a given wave vector $\mathbf{k}$. The initial condition then follows ~$\tilde{\mathbb{P}}(\mathbf{k},\vartheta,t=0|\vartheta_{0})=\delta(\vartheta-\vartheta_{0})$. Hence, $\tilde{\mathbb{P}}$ encodes the full information on the translational and angular motion.
@@
The formal solution of Sec. III is ~$\tilde{\mathbb{P}}=e^{(\mathcal{L}+\delta\mathcal{L}_{\mathbf{k}})t}\delta(\vartheta-\vartheta_{0})$ and to obtain explicit expressions a spectral-theory approach is applied on the operators. We shall also employ a time-dependent perturbation theory for small wave vectors $\mathbf{k}$ to generate the low-order moments. The following derivation is already outlined in Ref. Chepizhko and Franosch (2022) (see also Lapolla et al. (2020); Kurzthaler and Franosch (2017)) and for completeness, we summarize the most important steps.
@@
where we use the generalized angular basis $\{|\vartheta\rangle\}$ which is orthogonal in the following sense $\langle\vartheta|\vartheta_{0}\rangle=\delta(\vartheta-\vartheta_{0})$ and (over-) complete $\int_{0}^{2\pi}\mathrm{d}\vartheta|\vartheta\rangle\langle\vartheta|=\mathbb{1}$. We rely on the isomorphism between the periodic square-integrable functions $f(\vartheta),g(\vartheta)\in L^{2}[0,2\pi]$ and states $|f\rangle$,$|g\rangle$ in the Hilbert space $\mathcal{H}$. We are interested in the eigenfunctions $\{|\alpha\rangle:\alpha\in\mathbb{Z}\}$ of the unperturbed operator $\mathcal{L}$ for the standard orthonormal basis in $\mathcal{H}$, with the real-space representation $\langle\vartheta|\alpha\rangle=\exp(\textsf{i}\alpha\vartheta)/\sqrt{2\pi}$. The unperturbed operator $\mathcal{L}$ can be constructed in this basis by its matrix elements
@@
~$\tilde{\mathbb{P}}=\langle\vartheta|e^{(\mathcal{L}+\delta\mathcal{L}_{\mathbf{k}})t}\vartheta_{0}\rangle,$ (10)
@@ IV Harmonic Approximation
$\textsf{Var}[\mathbf{n}\cdot\langle\Delta\mathbf{r}(t)\rangle]=\langle[\mathbf{n}\cdot\Delta\mathbf{r}(t)-\mathbf{n}\cdot\langle\Delta\mathbf{r}(t)\rangle]^{2}\rangle$
@@
$\textsf{Var}[\mathbf{n}\cdot\langle\Delta\mathbf{r}(t)\rangle]=2D_{\mathbf{n}}(t-\tau+\tau e^{-t/\tau}),$ (23)
@@
$F(\mathbf{k},t)=\exp(-\textsf{i}k\mathbf{n}\cdot\langle\Delta\mathbf{r}(t)\rangle-\frac{k^{2}}{2}\textsf{Var}[\mathbf{n}\cdot\langle\Delta\mathbf{r}(t)\rangle]),$ (25)
@@ VII Exact Low-Order moments
For the case of a strong orienting torque $\gamma>\omega$, the prefactor of diffusive growth tproportional-$\propto t$ is much bigger for the perpendicular direction than for the parallel direction, i.e., the particle diffuses mainly in the direction perpendicular to the mean motion. In the harmonic approximation this becomes exact as the diffusion coefficient, $D_{\textbf{n}}=0$, evaluates to zero for the parallel direction, compare Eq. 23. The approximation coincides with the curves, the better, the stronger the orienting torque. For the case of $\gamma=1.05\omega$, the harmonic approximation does not match the analytic curve. The reason has been identified in Chepizhko and Franosch (2022) as the particle hops over the barrier before relaxing in the minimum.
@@ Appendix C DDM corrected with drift
Differential dynamic microscopy (DDM) has been first introduced by Cerbino and Trappe Cerbino and Trappe (2008) to measure the ISF experimentally, relying only on advanced imaging techniques, rather than scattering experiments. The theoretical background of DDM is nicely presented in Refs. Giavazzi et al. (2009); Wilson et al. (2011). This technique can also be adapted to the comoving frame, indicated in the literature as well Edera et al. (2021); Giavazzi and Cerbino (2014); Richards et al. (2021). In the DDM method, a sequence of microscopic images is taken, and dynamic contrast is measured. In the comoving frame, the image is shifted with velocity $v$ to include a drift and so the difference of the two images recorded at time $t$ and 00 is
@@ Appendix D Fourth-order moment
×$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle+O(|\mathbf{k}|^{5}).$ (50)
@@
×$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle=\frac{t^{4}}{24}\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle^{4}$
@@
×$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle$
@@
×$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle$
@@
×$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle$
@@
×$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle.$

### v2

@@ III Derivation of the intermediate scattering function
Besides the equations of motion for the individual trajectories from Sec. II, dynamical properties are equivalently encoded in the propagator $\mathbb{P}\coloneq\mathbb{P}(\mathbf{r},\vartheta,t|\vartheta_{0})$, which is defined as the conditional probability distribution of the particle displaced by $\mathbf{r}$ at lag time $t$ with orientation $\vartheta$, given it started with orientation $\vartheta_{0}$. By standard means Risken (1996), the associated Fokker-Planck equation can be derived
@@
It is favorable to solve directly for the spatial Fourier transform of the propagator $\tilde{\mathbb{P}}\coloneq\tilde{\mathbb{P}}(\mathbf{k},\vartheta,t|\vartheta_{0})$
$\tilde{\mathbb{P}}=\int_{\mathbb{R}^{2}}\!\mathrm{d}\,\textbf{r}\exp(-\textsf{i}\textbf{k}\cdot\textbf{r})\mathbb{P}(\mathbf{r},\vartheta,t|\vartheta_{0}).$ (5)
@@
$\eqcolon(\mathcal{L}+\delta\mathcal{L}_{\mathbf{k}})\tilde{\mathbb{P}}$, (6)
where the operator $\mathcal{L}$ encodes the angle-dependent motion containing the drifts $\omega$ and $\gamma$, and the rotational diffusion. The $\mathbf{k}$-dependent operator $\delta\mathcal{L}_{\mathbf{k}}$ encodes the deterministic active motion for a given wave vector $\mathbf{k}$. The initial condition then follows $\tilde{\mathbb{P}}(\mathbf{k},\vartheta,t=0|\vartheta_{0})=\delta(\vartheta-\vartheta_{0})$. Hence, $\tilde{\mathbb{P}}$ encodes the full information on the translational and angular motion.
@@
The formal solution of Sec. III is $\tilde{\mathbb{P}}=e^{(\mathcal{L}+\delta\mathcal{L}_{\mathbf{k}})t}\delta(\vartheta-\vartheta_{0})$ and to obtain explicit expressions a spectral-theory approach is applied on the operators. We shall also employ a time-dependent perturbation theory for small wave vectors $\mathbf{k}$ to generate the low-order moments. The following derivation is already outlined in Ref. Chepizhko and Franosch (2022) (see also Lapolla et al. (2020); Kurzthaler and Franosch (2017)) and for completeness, we summarize the most important steps.
@@
where we use the generalized angular basis $\{|\vartheta\rangle\}$ which is orthogonal in the following sense $\langle\vartheta|\vartheta_{0}\rangle=\delta(\vartheta-\vartheta_{0})$ and (over-) complete $\int_{0}^{2\pi}\mathrm{d}\vartheta|\vartheta\rangle\langle\vartheta|=\mathbb{1}$. We rely on the isomorphism between the periodic square-integrable functions $f(\vartheta),g(\vartheta)\in L^{2}[0,2\pi]$ and states $|f\rangle$$|g\rangle$ in the Hilbert space $\mathcal{H}$. We are interested in the eigenfunctions $\{|\alpha\rangle:\alpha\in\mathbb{Z}\}$ of the unperturbed operator $\mathcal{L}$ for the standard orthonormal basis in $\mathcal{H}$, with the real-space representation $\langle\vartheta|\alpha\rangle=\exp(\textsf{i}\alpha\vartheta)/\sqrt{2\pi}$. The unperturbed operator $\mathcal{L}$ can be constructed in this basis by its matrix elements
@@
$\tilde{\mathbb{P}}=\langle\vartheta|e^{(\mathcal{L}+\delta\mathcal{L}_{\mathbf{k}})t}\vartheta_{0}\rangle,$ (10)
@@ IV Harmonic Approximation
$\textsf{Var}[\mathbf{n}\cdot\Delta\mathbf{r}(t)]=\langle[\mathbf{n}\cdot\Delta\mathbf{r}(t)-\mathbf{n}\cdot\langle\Delta\mathbf{r}(t)\rangle]^{2}\rangle$
@@
$\textsf{Var}[\mathbf{n}\cdot\Delta\mathbf{r}(t)]=2D_{\mathbf{n}}(t-\tau+\tau e^{-t/\tau}),$ (23)
@@
$F(\mathbf{k},t)=\exp(-\textsf{i}k\mathbf{n}\cdot\langle\Delta\mathbf{r}(t)\rangle-\frac{k^{2}}{2}\textsf{Var}[\mathbf{n}\cdot\Delta\mathbf{r}(t)]),$ (25)
@@ VII Exact Low-Order moments
For the case of a strong orienting torque $\gamma>\omega$, the prefactor of diffusive growth $\propto t$ is much bigger for the perpendicular direction than for the parallel direction, i.e., the particle diffuses mainly in the direction perpendicular to the mean motion. In the harmonic approximation this becomes exact as the diffusion coefficient, $D_{\textbf{n}}=0$, evaluates to zero for the parallel direction, compare Eq. 23. The approximation coincides with the curves, the better, the stronger the orienting torque. For the case of $\gamma=1.05\omega$, the harmonic approximation does not match the analytic curve. The reason has been identified in Chepizhko and Franosch (2022) as the particle hops over the barrier before relaxing in the minimum.
@@ Appendix C DDM corrected with drift
Differential dynamic microscopy (DDM) has been first introduced by Cerbino and Trappe Cerbino and Trappe (2008) to measure the ISF experimentally, relying only on advanced imaging techniques, rather than scattering experiments. The theoretical background of DDM is nicely presented in Refs. Giavazzi et al. (2009); Wilson et al. (2011). This technique can also be adapted to the comoving frame, indicated in the literature as well Edera et al. (2021); Giavazzi and Cerbino (2014); Richards et al. (2021). In the DDM method, a sequence of microscopic images is taken, and dynamic contrast is measured. In the comoving frame, the image is shifted with velocity $v$ to include a drift and so the difference of the two images recorded at time $t$ and 0 is
@@ Appendix D Fourth-order moment
$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle+O(|\mathbf{k}|^{5}).$ (50)
@@
$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle=\frac{t^{4}}{24}\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle^{4}$
@@
$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle$
@@
$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle$
@@
$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle$
@@
$\times\langle l_{0}|\delta\mathcal{L}_{\mathbf{k}}r_{n}\rangle\langle l_{n}|\delta\mathcal{L}_{\mathbf{k}}r_{m}\rangle\langle l_{m}|\delta\mathcal{L}_{\mathbf{k}}r_{p}\rangle\langle l_{p}|\delta\mathcal{L}_{\mathbf{k}}r_{0}\rangle.$